OpenStudy (anonymous):

Solution of 2/x+1=x+1/2

OpenStudy (solomonzelman):

 cross multiply... can you do it, or need help cross multiplying ?

OpenStudy (anonymous):

No thanks I'm fine:)

OpenStudy (solomonzelman):

without simplifying the cross multiplication...\[(x+1)(x+1)=2 \times 2\]\[(x+1)^2=2^2\]\[x+1=2~~~~~~~~~~(\rm ~took~~ro ot~~of~~bo th~~sides)\]